# Review on the decoding algorithms for surface codes

### v4

@@ 5.3.3 Enhanced Belief Propagation: Quantum Ordered Statistics Decoding
OSD-$\mathbf{w}$: Higher order OSD is similar to OSD-0, the difference being that we now consider solutions $\mathbf{e}_{\Lambda_{w}}=[e_{[J]},e_{[\bar{J}]}]$ where $e_{[\bar{J}]}\neq\mathbf{0}$. OSD-$w$ begins by running the first five steps of OSD-0 and computing $e_{[J]}$, which for the sake of notation we will now denote by $e^{w=0}_{[J]}$. Once this is done, different candidates $\mathbf{e}_{k}$ can be found by making choices for $e_{[\bar{J}]}$ and solving
$\mathbf{e}_{\Lambda_{w}}=[e_{[J]},e_{[\bar{J}]}]=[e^{w=0}_{[J]}+\Lambda^{-1}_{[J]}\Lambda_{[\bar{J}]}e_{[\bar{J}]},\ e_{[\bar{J}]}],$ (18)
@@
where $\Lambda_{[\bar{J}]}$ is the submatrix obtained by taking the columns of $\Lambda$ indexed by $\bar{J}$. The solution $\mathbf{e}_{w}$ given in (18) satisfies the OSD syndrome equation
$\Lambda\mathbf{e}_{\Lambda_{w}}=s$ (19)
@@
for any choice of $e_{[\bar{J}]}$. The premise behind OSD-$w$ (considering $e_{[\bar{J}]}\neq\mathbf{0}$) is to find solutions $\mathbf{e}_{\Lambda_{w}}$ of lower Hamming weight than $\mathbf{e}_{\Lambda_{0}}$. The $e_{[\bar{J}]}\neq\mathbf{0}$ component has dimension $n-\text{rank}(H)$, implying that testing all possible configurations for $e_{[\bar{J}]}$ is intractable beyond a small value of $n$. For this reason, it is important to design a strategy that makes good choices for the $e_{[\bar{J}]}$ candidates. This is best approached [^25] A different strategy was employed in the works that first introduced qOSD. However, slight performance improvements were shown in [122] when using the combination sweep strategy. using the so called combination sweep strategy. This greedy search method works as follows:
@@
    At the start of the search, the OSD-$w$ solution is equal to the unordered OSD-0 solution: $\mathbf{e}_{\Lambda_{w}}=\mathbf{e}_{\Lambda_{0}}$.
@@
    Sort the bits in the $e_{[\bar{J}]}$ subvector of the solution according to the soft-outputs of the original BP decoding attempt. Note that this step is already built into OSD-0 when re-arranging the parity check matrix according to the BP outputs.
@@
    Test all possible weight-$1$ configurations of $e_{[\bar{J}]}$. There are a total of $n-\text{rank}(H)$ weight-11 candidates. If the weight of any of the candidates is lower than the weight of $\mathbf{e}_{\Lambda_{0}}$, update the choice of $\mathbf{e}_{\Lambda_{w}}$.
@@
    Try all possible weight-$2$ configurations in the first $w$ bits of $e_{[\bar{J}]}$. Obviously, the order [^26] It is worth mentioning that in [122], the order $w$ of the OSD algorithm is referred to as the search depth. $w$ will be upper bounded by $n-\text{rank}(H)$ the dimension of $e_{[\bar{J}]}$. The number of candidates is given by the binomial coefficient $\binom{w}{2}$. If the weight of any of the candidates is lower than the weight of the current choice for the solution, update the choice of $\mathbf{e}_{\Lambda_{w}}$.
@@
    The final step is analogous to that of the order 0 version: map the solution to the original bit-index order accordingly: $\mathbf{e}_{\Lambda_{w}}\rightarrow\mathbf{e}_{\text{OSD-w}}$.
OSD-$w$ entails testing a total of $n-\text{rank}(H)+\binom{w}{2}$ candidates for $\mathbf{e}_{\Lambda_{w}}$, out of which the minimum Hamming weight solution, or at least a better choice than $\mathbf{e}_{\Lambda_{0}}$, will have been found. As the order $w$ of the combination sweep strategy is increased, the likelihood of finding the solution of minimum Hamming weight to (19) increases. At the same time, this also implies additional computational demand, negatively impacting the complexity of the algorithm and its runtime performance.
@@ 5.3.4 OSD Complexity
In [46], the authors show that most of the performance improvements provided by qOSD postprocessing are achieved with OSD-0. While increasing the order $w$ of the algorithm yields benefit over setting $w=0$, for various QLDPC code families the improvement provided by running the algorithm with $w>0$ is marginal. It is important to mention that in [46], the authors use a different algorithm to the combination sweep strategy [122]. Instead, they apply an exhaustive approach where they test all possible permutations in the first $w$ bits of $e_{[\bar{J}]}$. Given that OSD-0 requires the solving of a linear system, its complexity will be at most $O(n^{3})$ (although there are matrix inversion algorithms with better complexity [155], they are quite impractical). The exhaustive approach to OSD-$w$ has a complexity in the general case of $O(n^{3}+n2^{w})$. Although the combination sweep also has an edge in terms of complexity ($\approx O(w^{\alpha}n^{3})$), it is likely that OSD-0 is the only version of the algorithm that can be successfully implemented in a real time system [122, 156]. Recently, an OSD-inspired reduced complexity approach known as stabilizer inactivation has been proposed. In [157] the authors showed that this strategy, which has $O(n^{2}\text{log}n)$ in the worst case, achieved a higher threshold for the family of generalized bicycle codes.
@@ 5.3.6 Measurement errors
There have been many proposals for handling circuit-level noise by using belief propagation. An important proposal for handling this type of errors in the surface code is the decoder deemed as belief-matching [87]. This decoder is base on a combination between the BP and MWPM algorithms, idea that was previously explored in [158] with perfect syndrome measurements. The original belief-matching algorithm achieved a $17.76\%$ threshold for the rotated planar code when depolarizing noise was considered over the data qubits. The decoding complexity of the algorithm was maintained when compared to MWPM when parallel processing was allowed [158]. Due to the impressive performance of this BP+MWPM approach, such method was generalized to handle noisy gates and SPAM errors in [87], resulting in one of the most vainglorious algorithms for obtaining logical qubits in such planar architecture codes. The basic idea behind the belief-matching algorithm is similar to the BPOSD decoder in the sense that the soft ouptuts of the sum-product algorithm are used for the posterior decoding algorithm to make use of them. In this sense, the weights of the the graph in the MWPM decoder are produced by using such a posteriori information. Regarding the generalization of the belief-matching for circuit-level noise, the authors discussed how to construct the circuit-level Tanner graph that takes the measurement circuits of the surface code into account so that belief propagation can be run over such graph and then use the soft output for obtaining the MWPM solution [87]. Note that $\mathcal{O}(d)$ measurement rounds are required for this. The authors found a threshold of $0.94\%$ for the belief-matching algorithms with circuit level noise, which is comparable to the belief-find decoder also proposed in such article, which follows the same procedure but using UF instead of MWPM and achieves a threshold of $0.937\%$ [87]. The authors discuss that their performances are similar due to the fact that most of the information needed for decoding is provided by the BP part of the algorithm.
@@
Another approach to deal with noisy syndromes is to consider soft syndrome information for the BP graph used for decoding as it was done for QLDPC codes in [159]. In such work, the authors discuss the fact that when a measure met is noisy, the input information of the Tanner graph regarding the information of the syndrome can be soft, i.e. a probability distribution conditional to the obtained noisy measurement. Therefore, the fact that the measurement outcome might not be precise is also fed to the BP algorithm. However, this study only considered the fact that the noisy measurements are a result of SPAM errors, neglecting circuit level noise. The authors discuss that this is considered to be future work in this direction. Interestingly, such paper was based on the previous work by Pattison et al. that discussed the use of soft measurement information in the context of MWPM and UF for surface codes [160]. The authors concluded that their modified decoders using soft information improved the threshold obtained by hard decision decoders for the circuit-level noise considered. In this sense, considering this approaches for BP decoders discussed for the surface code might be an interesting path to follow for dealing with circuit level noise.
@@
Finally, single-shot decoding using BPOSD was investigated for the 4D toric code in [161]. Single-shot decoding refers to estimating the error when noisy measurements are present in a single measurement round, i.e. without needing to measure the usual $\mathcal{O}(d)$ rounds. In such article, the authors propose to decode data qubit errors and noisy syndrome measurements altogether in a single stage via the BPOSD decoder. The authors discuss that they obtain better thresholds than using multiple-measurements and other single-shot decoder such as the cellular automaton decoder. However, they use a phenomenological noise model for faulty syndrome measurements as discussed before, i.e. they do not consider full circuit-level noise, which they consider to be future work [161]. Single-shot decoding is being actively investigated at the time of writing, specially from the point of view of QLDPC codes [162, 163, 164, 165]. This approach was originally proposed for some families of topological codes, and due to its advantage in terms of measurement rounds and, presumably, performance, seems to be interesting to study more codes that may admit this kind of decoding to deal with circuit-level noise.
@@ 5.4 The Tensor Network decoder
Tensor Network (TN) decoders are decoding methods, proposed by Bravyi et al. for the surface code, that aim to resolve the DQMLD problem, i.e. to estimate which is the most probable logical coset based on the syndrome information [121]. Note that, up to this point, all decoders considered have followed a QMLD logic, that is, their aim has been to seek for the most probable Pauli error (QMLD) since, as has been seen earlier, the error recovery problem contemplating error logical classes (DQMLD) belongs to the #P\#P complexity class [55]. For this reason, the TN decoder is also usually referred as the maximum likelihood decoder (MLD) [^27] Note that this nomenclature might be somewhat confusing as it does not explicitly represent the fact that it is a degenerated decoder.[121]. Fortunately, surface codes admit a natural representation in terms of Tensor Networks (TN) [166], feature that can be used for approximating the DQMLD problem targeting the most probable error coset. Decoding quantum error correcting codes with TNs was first considered in [167], where the authors described for the first time the equivalence between decoding a quantum code and contracting a TN for quantum turbo, polar and branching multiscale entanglement renormalization ansatz (MERA) codes [168, 169]. Later, the TN decoding was particularized for the planar code in [121], and, afterwards, in [170], the method was expanded to the rotated planar code yielding significant results. Lastly, in [171], the method was generalized to any quantum 2D code.
Planar codes are not the only class of topological codes that admit the TN decoding methods as, in general, those admit a natural representation in terms of 2D Projected Entangled Pair States (PEPS) [172]. For instance, the ground states of the Toric Code Hamiltonian, codifying two logical qubits, can be written exactly as PEPS with low bond dimension, and the same is true for all other surface codes such as quantum double models, color codes, and even string-net models [173]. This analytical correspondence implies that decoding such codes can be done using standard TN techniques, in particular those related to the contraction of 2D TNs. Such algorithms have been used widely in the calculation of partition functions of 2D classical lattice models [174, 175], as well as in the calculation of expectation values of 2D quantum states on a lattice [172, 176]. Take for instance the straightforward example of the Toric Code. The expectation value of loop operators around the torus geometry decode the value of the logical qubits, in a completely analogue way to the expectation value of Wilson loop operators in lattice gauge theories. As such, these loop operators are built from tensor products of Pauli matrices along the sites on a loop, and the TN contraction involved to estimate them can be done via usual contraction methods for 2D TNs.
@@ 5.4.1 Introduction to TNs
Figure 15: Some examples of tensor operations. a) indicates a scalar product of two vectors, b) indicates the trace of a matrix, c) indicates the product of a vector with a matrix and d) indicates a matrix product state.
@@
TNs accept an intuitive notation. FIG. 15 depicts some examples of tensors represented through various plots, where vectors are represented as circles, and matrices and higher order tensors, as squares. In a TN, some, or all the indices can be contracted according to a specific pattern, as shown in FIG. 15. If all the indices are contracted, the TN results in a scalar, as in the case of the two first examples of FIG. 15. The lines connecting tensors from one to the other represent the indices such that there is a sum over all their possible values, whereas lines that are not connecting to anywhere represent free indices. Those indices that correspond to the original physical Hilbert spaces are called physical indices, and the rest (connecting the different tensors with each other) are called bond indices, and are responsible for the entanglement in the quantum many-body wavefunction. The range of values of a bond index is called bond dimension, following common TN terminology [166, 121].
@@
Consider now the following case: we have a 2D tensor network such as the one showed in FIG. 16, which is the one encountered when computing expectation values of a projected entanglement pair state (PEPS). The constituent tensors on the bulk have bond dimension 4, while the ones on the boundary have lower bond dimensions. One can consider the left-most (or right-most) column as a matrix product state (MPS) and columns within the bulk of the PEPS as matrix product operators (MPO). When contracting the left MPS with its nearest MPO, the overall 2D TN will be reduced by one column as shown in FIG. 16, and for the new column which is obtained the vertical bond dimension will have increased in an exponential manner. Contracting the entire 2D TN will most likely be an unfeasible challenge as a result of the exponentially increasing bond dimension. Fortunately, we can still obtain meaningful results for continuous column contraction while avoiding an exponential growth [177, 178]. For the cases we consider, this is made by using the QR matrix decomposition altogether with the singular value decomposition (SVD), which is used to truncate the vertical bond-dimensions [121, 179]. Truncating the bond dimension of the column contractions will yield an approximate result to the overall contraction as opposed to an analytical result which, nevertheless, allows for a defined run time since, when the contraction results in a last column, its vertical contraction can be carried at polynomial time [171].
@@
Figure 16: A tensor network which contracts the left most column with the second one.

### v5

@@ 5.3.3 Enhanced Belief Propagation: Quantum Ordered Statistics Decoding
OSD-$\mathbf{w}$: Higher order OSD is similar to OSD-0, the difference being that we now consider solutions $\hat{e}_{\Lambda_{w}}=[\hat{e}_{[J]},\hat{e}_{[\bar{J}]}]$ where $\hat{e}_{[\bar{J}]}\neq\mathbf{0}$. OSD-$w$ begins by running the first five steps of OSD-0 and computing $\hat{e}_{[J]}$, which for the sake of notation we will now denote by $\hat{e}^{w=0}_{[J]}$. Once this is done, different candidates $\mathbf{\hat{e}}_{k}$ can be found by making choices for $\hat{e}_{[\bar{J}]}$ and solving
$\hat{e}_{\Lambda_{w}}=[\hat{e}_{[J]},\hat{e}_{[\bar{J}]}]=[\hat{e}^{w=0}_{[J]}+\Lambda^{-1}_{[J]}\Lambda_{[\bar{J}]}\hat{e}_{[\bar{J}]},\ \hat{e}_{[\bar{J}]}],$ (27)
@@
where $\Lambda_{[\bar{J}]}$ is the submatrix obtained by taking the columns of $\Lambda$ indexed by $\bar{J}$. The solution $\hat{e}_{w}$ given in (27) satisfies the OSD syndrome equation
$\Lambda\hat{e}_{\Lambda_{w}}=\bar{s}$ (28)
@@
for any choice of $\hat{e}_{[\bar{J}]}$. The premise behind OSD-$w$ (considering $\hat{e}_{[\bar{J}]}\neq\mathbf{0}$) is to find solutions $\mathbf{e}_{\Lambda_{w}}$ of lower weight than $\hat{e}_{\Lambda_{0}}$. The $\hat{e}_{[\bar{J}]}\neq\mathbf{0}$ component has dimension $n-\text{rank}(H)$, implying that testing all possible configurations for $\hat{e}_{[\bar{J}]}$ is intractable beyond a small value of $n$. For this reason, it is important to design a strategy that makes good choices for the $\hat{e}_{[\bar{J}]}$ candidates. This is best approached [^31] A different strategy was employed in the works that first introduced qOSD. However, slight performance improvements were shown in [98] when using the combination sweep strategy. using the so called combination sweep strategy. This greedy search method works as follows:
@@
    At the start of the search, the OSD-$w$ solution is equal to the unordered OSD-0 solution: $\hat{e}_{\Lambda_{w}}=\hat{e}_{\Lambda_{0}}$.
@@
    Sort the bits in the $\hat{e}_{[\bar{J}]}$ subvector of the solution according to the soft-outputs of the original BP decoding attempt. Note that this step is already built into OSD-0 when re-arranging the parity check matrix according to the BP outputs.
@@
    Test all possible weight-$1$ configurations of $\hat{e}_{[\bar{J}]}$. There are a total of $n-\text{rank}(H)$ weight-11 candidates. If the weight of any of the candidates is lower than the weight of $\hat{e}_{\Lambda_{0}}$, update the choice of $\hat{e}_{\Lambda_{w}}$.
@@
    Try all possible weight-$2$ configurations in the first $w$ bits of $\hat{e}_{[\bar{J}]}$. Obviously, the order [^32] It is worth mentioning that in [98], the order $w$ of the OSD algorithm is referred to as the search depth. $w$ will be upper bounded by $n-\text{rank}(H)$ the dimension of $\hat{e}_{[\bar{J}]}$. The number of candidates is given by the binomial coefficient $\binom{w}{2}$. If the weight of any of the candidates is lower than the weight of the current choice for the solution, update the choice of $\hat{e}_{\Lambda_{w}}$.
@@
    The final step is analogous to that of the order 0 version: map the solution to the original bit-index order accordingly: $\hat{e}_{\Lambda_{w}}\rightarrow\hat{e}_{\text{OSD-w}}$.
OSD-$w$ entails testing a total of $n-\text{rank}(H)+\binom{w}{2}$ candidates for $\hat{e}_{\Lambda_{w}}$, out of which the minimum Hamming weight solution, or at least a better choice than $\hat{e}_{\Lambda_{0}}$, will have been found. As the order $w$ of the combination sweep strategy is increased, the likelihood of finding the solution of minimum Hamming weight to (28) increases. At the same time, this also implies additional computational demand, negatively impacting the complexity of the algorithm and its runtime performance.
@@ 5.3.4 OSD Complexity
The complexity of the BP process is $\mathcal{O}(nj)$ [129, 130], where $n$ indicates the number of physical qubits and $j$ is the average number of non-trivial elements per column in the parity check matrix. Moreover, in [37], the authors show that most of the performance improvements provided by qOSD postprocessing are achieved with OSD-0. While increasing the order $w$ of the algorithm yields benefit over setting $w=0$, for various QLDPC code families the improvement provided by running the algorithm with $w>0$ is marginal. It is important to mention that in [37], the authors use a different algorithm to the combination sweep strategy [98]. Instead, they apply an exhaustive approach where they test all possible permutations in the first $w$ bits of $e_{[\bar{J}]}$. Given that OSD-0 requires the solving of a linear system, its complexity will be at most $\mathcal{O}(n^{3})$ (although there are matrix inversion algorithms with better complexity, they are quite impractical). The exhaustive approach to OSD-$w$ has a complexity in the general case of $\mathcal{O}(n^{3}+n2^{w})$. Although the combination sweep also has an edge in terms of complexity ($\approx\mathcal{O}(w^{\alpha}n^{3})$), it is likely that OSD-0 is the only version of the algorithm that can be successfully implemented in a real time system [98, 131]. Recently, an OSD-inspired reduced complexity approach known as stabilizer inactivation has been proposed. In [132] the authors showed that this strategy, which has $\mathcal{O}(n^{2}\text{log}n)$ in the worst case, achieved a higher threshold for the family of generalized bicycle codes.
@@ 5.3.6 Measurement errors
There have been many proposals for handling circuit-level noise by using belief propagation. An important proposal for handling this type of errors in the surface code is the decoder deemed as belief-matching [65]. This decoder is base on a combination between the BP and MWPM algorithms, idea that was previously explored in [133] with perfect syndrome measurements. The original belief-matching algorithm achieved a $17.76\%$ threshold for the rotated planar code when depolarizing noise was considered over the data qubits. The decoding complexity of the algorithm was maintained when compared to MWPM when parallel processing was allowed [133]. Given the impressive performance of this BP+MWPM approach, it was generalized to handle noisy gates and SPAM errors in [65], resulting in an excellent decoding strategy for such planar codes. The basic idea behind the belief-matching algorithm is similar to the BPOSD decoder in the sense that the soft ouptuts of the sum-product algorithm are used for the posterior decoding algorithm to make use of them. In this sense, the weights of the the graph in the MWPM decoder are produced by using such a posteriori information. Regarding the generalization of the belief-matching for circuit-level noise, the authors discussed how to construct the circuit-level Tanner graph that takes the measurement circuits of the surface code into account so that belief propagation can be run over such graph and then use the soft output for obtaining the MWPM solution [65]. Note that $\mathcal{O}(d)$ measurement rounds are required for this. The authors found a threshold of $0.94\%$ for the belief-matching algorithms with circuit level noise, which is comparable to the belief-find decoder also proposed in such article, which follows the same procedure but using UF instead of MWPM and achieves a threshold of $0.937\%$ [65]. The authors discuss that their performances are similar due to the fact that most of the information needed for decoding is provided by the BP part of the algorithm.
@@
Another approach to deal with noisy syndromes is to consider soft syndrome information for the BP graph used for decoding as it was done for QLDPC codes in [134]. In such work, the authors discuss the fact that when a measurement is noisy, the input information of the Tanner graph regarding the information of the syndrome can be soft, i.e. a probability distribution conditional to the obtained noisy measurement. Therefore, the fact that the measurement outcome might not be precise is also fed to the BP algorithm. However, this study only considered the fact that the noisy measurements are a result of SPAM errors, neglecting circuit level noise. The authors discuss that this is considered to be future work in this direction. Interestingly, such paper was based on the previous work by Pattison et al. that discussed the use of soft measurement information in the context of MWPM and UF for surface codes [135]. The authors concluded that their modified decoders using soft information improved the threshold obtained by hard decision decoders for the circuit-level noise considered. In this sense, considering this approaches for BP decoders discussed for the surface code might be an interesting path to follow for dealing with circuit level noise.
Moreover, for the latest articles of [123, 124], an approach was taken where, via the software package [111], one can compute the parity check matrix of the circuit-level noise, which contains all single error mechanisms in the syndrome extraction circuit as columns. Once this parity check matrix is obtained, given a syndrome, one can return an error applying BP+OSD. Applying BPOSD to circuit-level noise yields very good results although it is under constraint of considering a much larger parity check matrix to be inverted. Refer to Appendix B for a more detailed description of the multiple measurement-round measurement extraction and decoding.
@@
Finally, single-shot decoding using BPOSD was investigated for the 4D toric code in [136]. Single-shot decoding refers to estimating the error when noisy measurements are present in a single measurement round, i.e. without needing to measure the usual $\mathcal{O}(d)$ rounds. In such article, the authors propose to decode data qubit errors and noisy syndrome measurements altogether in a single stage via the BPOSD decoder. The authors discuss that they obtain better thresholds than using multiple-measurements and other single-shot decoder such as the cellular automaton decoder. However, they use a phenomenological noise model for faulty syndrome measurements as discussed before, i.e. they do not consider full circuit-level noise, which they consider to be future work [136]. Single-shot decoding is being actively investigated at the time of writing, specially from the point of view of QLDPC codes [137, 138]. This approach was originally proposed for some families of topological codes, and due to its advantage in terms of measurement rounds and, presumably, performance, seems to be interesting to study more codes that may admit this kind of decoding to deal with circuit-level noise.
@@ 5.4 The Tensor Network decoder
Tensor Network (TN) decoders are decoding methods, proposed by Bravyi et al. for the planar code, that aim to resolve the DQMLD problem. For this reason, the TN decoder is also usually referred as the maximum likelihood decoder (MLD) [^33] Note that this nomenclature might be somewhat confusing as it does not explicitly represent the fact that it is a degenerated decoder.[60]. Fortunately, surface codes admit a natural representation in terms of Tensor Networks (TN) [139], feature that can be used for approximating the DQMLD problem targeting the most probable error coset. Decoding quantum error correcting codes with TNs was first considered in [140] . Later, the TN decoding was particularized for the planar code in [60], and, afterwards, in [141], the method was expanded to the rotated planar code yielding significant results. Lastly, in [142], the method was generalized to any quantum 2D code.
@@ 5.4.1 Introduction to TNs
Figure 16: Some examples of tensor operations. a) indicates a scalar product of two vectors as shown in equation (29), b) indicates the trace of a matrix as shown in equation (30), c) indicates the product of a vector with a matrix as shown in equation (31) and d) indicates a matrix product state as shown in equation (32).
TNs accept an intuitive notation. FIG. 16 depicts some examples of tensors represented through various plots, where vectors are represented as circles, and matrices and higher order tensors, as squares. Each plot represents the following expressions from left to right:
$\sum_{i}v_{i}w_{i},$ (29)
$\operatorname{Tr}(A),$ (30)
$\sum_{i}A_{i,j,k}v_{i},$ (31)
$\sum_{i,j}A_{k,m,i}B_{i,n,j}C_{j,o,l}.$ (32)
@@
In a TN, some, or all the indices can be contracted according to a specific pattern, as shown in FIG. 16. If all the indices are contracted, the TN results in a scalar, as in the case of the two first examples of FIG. 16. The lines connecting tensors from one to the other represent the indices such that there is a sum over all their possible values, whereas lines that are not connecting to anywhere represent free indices. Those indices that correspond to the original physical Hilbert spaces are called physical indices, and the rest (connecting the different tensors with each other) are called bond indices, and are responsible for the entanglement in the quantum many-body wavefunction. The range of values of a bond index is called bond dimension, following common TN terminology [60, 139].
@@
Consider now the following case: we have a 2D tensor network such as the one showed in FIG. 17, which is the one encountered when computing expectation values of a projected entanglement pair state (PEPS). The constituent tensors on the bulk have bond dimension 4, while the ones on the boundary have lower bond dimensions. One can consider the left-most (or right-most) column as a matrix product state (MPS) and columns within the bulk of the PEPS as matrix product operators (MPO). When contracting the left MPS with its nearest MPO, the overall 2D TN will be reduced by one column as shown in FIG. 17, and for the new column which is obtained, the vertical bond dimension will have increased in an exponential manner. Contracting the entire 2D TN will most likely be an unfeasible challenge as a result of the exponentially increasing bond dimension. Fortunately, we can still obtain meaningful results for continuous column contraction while avoiding an exponential growth [139]. For the cases we consider, this is made by using the QR matrix decomposition altogether with the singular value decomposition (SVD), which is used to truncate the vertical bond-dimensions [60]. Truncating the bond dimension of the column contractions will yield an approximate result to the overall contraction as opposed to an analytical result which, nevertheless, allows for a defined run time since, when the contraction results in a last column, its vertical contraction can be carried at polynomial time [142].
@@
Figure 17: A tensor network which contracts the left most column with the second one. When doing the contraction, the vertical bond dimensions are truncated to avoid exponential increase of the complexity.